OpenStudy (anonymous):

using implicit differentiation, find dy/dx of (x^2)y+x(y^2) = 6

OpenStudy (hba):

The easiest way to do this is to use implicit differentiation on it as it stands. You differentiate using the normal rules but you treat y as an unknown function of x. This means f(y) differentates to f '(y)*(dy/dx) by the chain rule. xy + x + y^2 = 6 First implicit differentiation gives [1*y + x*1*(dy/dx)] + 1 + 2y*(dy/dx) = 0 .....[from xy].....................[from y^2] You can rearrange this to give dy/dx = (function of x and y) Second implicit differentiation gives dy/dx + [1*(dy/dx) + x*(d2y/dx2)] + [2*(dy/dx)*(dy/dx) + 2y*(d2y/dx2)] = 0 ...................[from x*(dy/dx)]......................[from 2y*(dy/dx)] Now you replace each dy/dx by its function of x and y found above. Then rearrange it to give d2y/dx2 = (function of x and y).

OpenStudy (anonymous):

i don't understand how you split it up, i'm confused.... that honestly made no sense

OpenStudy (anonymous):

yay now my head hurts even worse

OpenStudy (hba):

How Much More Questions Do You Have ?

OpenStudy (hba):

Or Is This The Last Question ?

OpenStudy (anonymous):

this is my first question...

OpenStudy (hba):

How Does Your Head Hurt ?

OpenStudy (anonymous):

i've got about 16 more for this section, then 5 on a second section on it and a worksheet

OpenStudy (anonymous):

i'm a junior in highschool taking calculus.. this isn't making sense

OpenStudy (hba):

Oh Wow Genius :D

OpenStudy (anonymous):

obviously not if this isn't making sense. sorry it's just frustrating!

OpenStudy (anonymous):

and it's not making a bit of sense

OpenStudy (hba):

Well, (x^2)y + x(y^2) = 6 d/dx[(x^2)y + x(y^2)] = d/dx (6)

OpenStudy (hba):

2xy + dy/dx(x^2) + y^2 + 2y(dy/dx)x =0

OpenStudy (anonymous):

i'm lost stop

OpenStudy (anonymous):

how'd you get 2xy + dy/dx(x^2 + y^2 + 2y(dydx)x=0?

OpenStudy (hba):

You dont know how to differentiate ?

OpenStudy (anonymous):

I do, but it's confusing me

OpenStudy (anonymous):

the d/dx of x^2y?

OpenStudy (anonymous):

you use the multiplication rule right? udv + vdu?

OpenStudy (hba):

\[x ^{2}y+xy^2=6\]

OpenStudy (anonymous):

x^2(1) + y(2x) = 2xy + x^2

OpenStudy (hba):

Yes I am Using Product Rule

OpenStudy (anonymous):

then how did you get 2xy then dy/dx x^2?

OpenStudy (anonymous):

or do you get dy/dx and all that stuff in parenthasis because you're finding the d/dx of y

OpenStudy (hba):

\[\frac{ dy }{ dx }[(x^2)y + x(y^2)] =\frac{ dy }{ dx }6\]

OpenStudy (hba):

Now Use Product Rule

OpenStudy (hba):

\[2xy + dy/dx(x^2) + y^2 + 2y(dy/dx)x =0\]

OpenStudy (hba):

Now Take The Terms Including dy/dx On One Side

OpenStudy (hba):

\[dy/dx(x^2)+2y(dy/dx)x = -y^2-2xy\]

OpenStudy (hba):

Now Take Common dy/dx

OpenStudy (hba):

\[dy/dx(x^2+2xy)=-y^2-2xy\]

OpenStudy (anonymous):

then divide?

OpenStudy (hba):

So,\[\frac{ dy }{ dx }=\frac{ -(y^2+2xy)}{x^2+2xy }\]

OpenStudy (hba):

^ Now Thats Your Answer Got It :)

OpenStudy (anonymous):

question, when using the product rule

OpenStudy (anonymous):

how do you know which one to put dy/dx with

OpenStudy (anonymous):

like when we had x^2y and got dy/dx(x^2) + 2xy

OpenStudy (anonymous):

why'd we put dy/dx with x^2 and not 2xy?

OpenStudy (hba):

Well Its Your Own Decision Whether to Put U first or V First If We talk about the formula

OpenStudy (hba):



OpenStudy (anonymous):

okay but then which one do i put the dy/dx with?

OpenStudy (anonymous):

like the next problem is y2 = (x+1)/(x-1)

OpenStudy (anonymous):

so would i do 2y (dy/dx) = vdu - udv/v^2

OpenStudy (hba):

Depends Completely On You

OpenStudy (anonymous):

so dy/dx2y = (x+1)(1) - (x-1)(1) / (x-1)^2

OpenStudy (hba):

@Godsgirl Please Help Her Out I Gotta Leave- Thanks :)

OpenStudy (anonymous):

Hey so wht do u need help with

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

I'm here. I have a question on where to put the dy/dx when solving

OpenStudy (anonymous):

right now i'm working on y^2 = x-1/x+1

OpenStudy (anonymous):

so i take dy/dx y^2 = (x+1)(1) - (x-1)(1) / (x+1)^2 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and the left side will equal 2y when i take dy/dx of it

OpenStudy (anonymous):

now where do i keep dy/dx, when solving?

OpenStudy (anonymous):

here i already solved it

OpenStudy (anonymous):

where do i put it? because i know i need to get it alone

OpenStudy (anonymous):

I sent it to u on ur messages

OpenStudy (anonymous):

that's not the question i'm asking about..

OpenStudy (anonymous):

the new one is

OpenStudy (anonymous):

which one are u asking about then

OpenStudy (anonymous):

y^2 = x-1/x+1

OpenStudy (anonymous):

i'm not trying to come off as rude, i'm sorry!

OpenStudy (anonymous):

just frustrated

OpenStudy (hba):

I am Back

OpenStudy (hba):

\[y^2 = (x-1) /(x+1)\]

OpenStudy (hba):

differentiate both sides with respect to x

OpenStudy (anonymous):

oh okay i was lost there

OpenStudy (anonymous):

i see where shes at now

OpenStudy (hba):

\[2y dy/dx = [(x+1)-(x-1)] /(x+1)^2 \]

OpenStudy (hba):

\[2y dy/dx = 2/(x+1)^2\]

OpenStudy (hba):

\[dy/dx = y / (x+1)^2\]

OpenStudy (hba):

@jamroz Got It :)

OpenStudy (anonymous):

so you keep dy/dx with the y on the left

OpenStudy (anonymous):

but why? cause i had gotten that far, i just didn't know which side to leave dy/dx with?

OpenStudy (hba):

Now You Got Which Side You Had To Keep It :) Anyways I am Off To Sleep, Aww My Head Hurts Anyways Bye :)

OpenStudy (anonymous):

thank you so much!